-Kor Cor The Henrys law constant of oxygen in water at 25 °C is 773 atm mol kg of water. Calculate the molality of oxygen

NUMBER 10 PLEASE
Answer #1

8) ..Griviem..Kpa.(partial pressure 04.02.) – 0..20. atm... OR=Kpa..20. kPa... .. Henrys law constant for 2 - 773(atm / mall liassuming the same solubility in blood, this is very low amount of oxygen carried by 1 litre of blood .So 5 litre of blood carry 1.290 mMol oxygen .

Now considering the total Hb in 1 litre of blood then ..

13.5 to 17.5 gm/ dl of blood carry oxygen = 1.34 (ml of O2)*17.5

=23.45 ml of oxygen carry's by 1 DL of HB

Then 1 litre of blood carrys = 23.45*106

= 234.5*105 ml of oxygen

So with this much amount of oxygen the solubility is much low so without haemoglobin, the survival is too difficult without haemoglobin.